OpenStudy (civicsiscool44):

Parallelogram FGHI on the coordinate plane below represents the drawing of a horse trail through a local park: Quadrilateral FGHI is shown on the coordinate plane with coordinates F at negative 6, 6. G at negative 2, 6. H at negative 4, 3. and I negative 8, 3. Point A is at 0, 4 and point D is at negative 6, negative 5. In order to build a scale model of the trail, the drawing is enlarged on the coordinate plane. If two corners of the trail are at point A(0, 4) and point D(−6, −5), what is another point that could represent a corner of the trail? (9, −5) (6, −5) (12, −5) (3, −5

OpenStudy (civicsiscool44):

@flcon04

OpenStudy (flcon04):

Okay so this is a multi choice as well. So tell me whats your thoughts on this question?

OpenStudy (civicsiscool44):

I'm thinking it might be B or C but I'm not sure

OpenStudy (flcon04):

Okay what I'll do is do solve the answer myself but show the steps I took.

OpenStudy (flcon04):

hmmm i just solved it but its not a answer choice huh

OpenStudy (flcon04):

Ah!

OpenStudy (flcon04):

I almost positive its C! But my digit is 1 number off.

OpenStudy (flcon04):

Is there something missing from the question maybe?

OpenStudy (flcon04):

Nevermind its B! Im 100% positive this time. my bad i scaled wrong

OpenStudy (flcon04):

so here is my work, If there is the 2 points i can work out the base length by x2 - x1 7 - 3 = 4 so side a of the triangle above= 4 then side b y2 - y1 3 - 0 =3 so side b = 3 so h^2 = a^2 + b^2 h^2 = 4^2 + 3^2 h^2 = 16 + 9 h^2 = 25 so h = 5

OpenStudy (flcon04):

and the other side is 6 i was getting 12 last time sorry about that